# Spin precession effects in the phasing formula of eccentric compact binary inspirals till the second post-Newtonian order

### v1

@@ Abstract
Compact binary systems that emit gravitational waves (GW) are expected to exhibit orbital eccentricity together with generic spin configurations, leading to the precession of the orbital angular momentum, the individual spins, and the orbital plane around the total angular momentum. Although eccentric binaries with aligned spins have been extensively studied, closed-form post-Newtonian (PN) expressions that simultaneously capture both eccentricity and precessing-spin effects have remained unavailable. The presence of eccentricity significantly complicates orbital evolution, since solving the coupled differential equations generally requires numerical integration, thereby slowing waveform generation. In this work, we exploit the separation of timescales between the orbital motion, spin precession, and radiation reaction, and employ the precession-averaging method introduced by Morras et al. (2025) to effectively “average away” the explicit time dependence from the spin–orbit and spin–spin dynamics till the second PN order. Building on this framework, we derive analytic phasing formulae from the evolution equations for the orbital frequency and eccentricity, treating eccentricity as a small parameter. Closed-form solutions for the eccentricity evolution and gravitational-wave phase are obtained up to the eighth power of the initial eccentricity $e_{0}$. The TaylorT2 approximant is generalized to incorporate spin-precession effects, and the orbital phase is calculated in both the time and frequency domains. To further enhance applicability, a resummation is performed on the TaylorT2 phasing formula to improve accuracy for moderate to high initial eccentricities. Taken together, these results provide efficient, closed-form phasing expressions that consistently capture the interplay of eccentricity and precessing spins, thereby enabling accurate and computationally tractable GW modeling for data analysis.
@@ I Introduction
A key part that comes after parameter estimation is the inference of formation channels of neutron stars and stellar-mass black holes. Eccentricity [21, 22, 23, 24, 25, 26, 27] and spin-precession [28, 29, 30], among other observables, are extremely useful for tracing the formation channels of such compact objects. It is to be noted that the absence of either eccentricity or spin-precession from template banks can lead to significant biases in parameter estimation [31, 32, 33, 34, 35, 36]. Also, since both introduce amplitude modulations in the GW signal, it becomes important to introduce both effects in waveform models in order to un-entangle the degeneracy caused by the presence of the two. Furthermore, inclusion of all known physical effects in waveform models will help in performing precision tests of General Relativity (GR) [37, 38, 39, 40, 41, 42, 43, 44].
So far, there has been significant progress in the development of eccentric waveform models. One of the first inspiral-merger-ringdown (IMR) models incorporating eccentricity was ENIGMA [45, 46], followed recently by its spinning counterpart ESIGMA [47]. PN inspired or inspiral only models include [48, 49, 50, 51, 52], effective-one-body (EOB) models include [53, 54, 55, 56], whereas NR-informed models include [57]. Similarly, for precessing spins, PN calculations include [58, 59, 60, 61, 62], EOB models include [63, 64], Phenom models include [65, 66], and NR-informed models include [67]. However, not much progress has been made in the combined modeling of precession and eccentricity [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78].
@@
This paper is an update to both [48] and [90]. Compared to the aligned-spin eccentric phasing of [90], in this paper we generalize the phasing formula to precessing-spins in elliptical orbits. We give the phasing formula in both time and frequency domain till 2 PN, and till an accuracy of the eighth power of the initial eccentricity $e_{0}$. We use inputs from [77] to calculate two Taylor approximants: TaylorT2 and TaylorF2 in the time and frequency domain respectively. We find that the phasing formulae are functions of intrinsic orbital parameters like mass ratios, spins, and deformability factors.
@@
This paper is organized as follows: In Sec. II, we first present the inputs, that is, the evolution equations for the eccentricity weighted orbital frequency (or the PN parameter $y$) and the time-eccentricity $e_{t}$ (to be called as eccentricity henceforth), which we denote as $e$ in this paper. We then turn the evolution equations with respect to time as a single evolution equation for the eccentricity $e$ as a function of the PN parameter $y$ using chain rule. This particular equation can only be solved exactly at the Newtonian order for arbitrary eccentricities. However, for higher PN orders, the evolution equation simply becomes untractable by currently available methods. Nevertheless, as shown by [48], if one treats eccentricity as a small parameter, the differential equation (DE) can be solved exactly, and a closed-form solution can be found till some order of the initial eccentricity $e_{0}$. However, the coefficient of each PN order in the eccentricity expanded DE was constant for the non-spinning and even in the aligned spins case. For generic spins, the coefficients themselves are time dependent, leading to further complications in solving the eccentricity evolution DE. To get around it, Refs. [71, 77] showed that the early inspiral binary problem could be treated as a multiple scale analysis (MSA) problem, which is possible since timescales for the orbit, spin-precession, and radiation-reaction can be clearly segregated. This allows the PN coefficients dependent on the spins to be precession-averaged, effectively taking away the time-dependence from the spin dynamics at the radiation-reaction timescale. Hence, these effective constants can now be put back into the eccentricity evolution DE and then solved as an ordinary DE, which we perform. The eccentricity solution as a function of the eccentricity-weighted orbital frequency $y$ is then utilized to find the time to coalescence $t$ as a function of $y$, and consequently, the time-domain orbital phasing $\phi$, or TaylorT2. We then perform a simple resummation of the TaylorT2 phasing. Finally, we use the Shifted Uniform Asymptotics (SUA) method introduced by [62] to find the frequency domain phasing $\Psi$.
@@
In Sec. III, we first calculate the number of cycles a real GW signal sweeps across a particular detector frequency range. We compare three different spin scenarios (almost aligned, almost anti-aligned, and perpendicular) for three different detector configurations and mass ranges. This we show for two different initial eccentricities. We then calculate the number of cycles swept by the numerically calculated TaylorT2 phase from the evolution equations and compare it with the analytical small-eccentricity expanded, as well as the resummed TaylorT2 phase number of cycles. Through this comparison, we show the domain of validity of our analytic TaylorT2 phase. We show that the $e_{0}^{8}$ expanded TaylorT2 phase incurs less than one cycle, compared to the numerically evaluated TaylorT2 phase, for initial eccentricity $e_{0}$ as high as 0.7. We also show that the resummed TaylorT2 is valid for initial eccentricities as high as 0.8.
@@
where $\bm{S}_{i}=m_{i}^{2}\bm{\chi}_{i}$, and $\bm{\chi}_{A/S}=\frac{\bm{\chi}_{1}\mp\bm{\chi}_{2}}{2}$. However, the approximants in the supplementary material have been expressed in ($\bm{\chi}_{1},\bm{\chi}_{2},q_{1},q_{2}$) notation for ease of use for future calculations. We also use symmetric/anti-symmetric spin induced quadrupole moment constants as $\kappa_{A/S}=\frac{q_{1}\mp q_{2}}{2}$ in our results, where $q_{i}$ were defined in [77] after Eq. (C6). Furthermore, we replace all the individual masses $m_{i}$ with $\nu=\frac{m_{1}\,m_{2}}{M^{2}}$ and $\delta=\sqrt{1-4\nu}$, where $M$ is the total mass of the binary system.
@@ II.1 Eccentricity and PN parameter evolution equations
The evolution equations of $y$ and $e$ are given till 2 PN for fully spinning and eccentric orbits by
@@
where $\mathcal{D}\equiv\frac{M}{\left(1-e^{2}\right)^{3/2}}\frac{d}{dt}$, $y=\frac{\left(M\omega\right)^{1/3}}{\sqrt{1-e^{2}}}$, and the non-spinning (NS) coefficients are given till 2 PN by [77].
@@
The spin-orbit (SO) terms are given till 2 PN in terms of $\chi_{eff}=\bm{\hat{L}}\cdot\bm{s}$ and $\delta\chi=\bm{\hat{L}}\cdot\left(\bm{s}_{1}-\bm{s}_{2}\right)$ by [77]
@@ II.2 Eccentricity evolution
Eqs. (2) can be combined to obtain the eccentricity evolution with respect to the PN parameter $y$. The DE cannot be solved for arbitrary eccentricities using current methods and hence needs to be expanded in small eccentricity in order to obtain an analytical solution. In the main paper, we show the $e^{2}$ expanded DE. However, the $e^{8}$ expanded DE is displayed in a supplementary file. The $e^{2}$ expanded DE has the following form, after precession averaging of the various PN coefficients appearing in the evolution equations (2)
@@
$c_{\rm SS,circ}$ $=\left(\frac{5\nu}{24}-\frac{5}{48}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}-\frac{5\delta}{24}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)+\left(\frac{5}{48}-\frac{5\nu}{24}\right)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}+\frac{5\delta}{24}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}$
$+\left(\frac{5\nu}{24}-\frac{5}{48}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}+\left(\frac{5}{48}-\frac{5\nu}{24}\right)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\,$, (10a)
$c_{\rm NS,ecc}$ $=-\frac{19}{3y}+y\left(\frac{197\nu}{18}-\frac{2833}{504}\right)+y^{3}\left(\frac{833\nu^{2}}{216}-\frac{32537\nu}{1008}+\frac{26464169}{1524096}\right)\,$, (10b)
$c_{\rm SO,ecc}$ $=\frac{157\delta}{18}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}+\left(\frac{157}{18}-\frac{55\nu}{9}\right)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}-\frac{377\pi}{24}\,$, (10c)
$c_{\rm SS,ecc}$ $=\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)\left[\left(\frac{89\nu}{2}-\frac{89}{4}\right)\kappa_{A}-\frac{137\delta}{192}-\frac{89\delta\kappa_{S}}{4}\right]+\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}\left[-\frac{89\delta\kappa_{A}}{8}-\frac{2503\nu}{192}\right$.
$\left.+\left(\frac{89\nu}{4}-\frac{89}{8}\right)\kappa_{S}-\frac{137}{384}\right]+\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}\left[-\frac{89\delta\kappa_{A}}{8}+\frac{2777\nu}{192}+\left(\frac{89\nu}{4}-\frac{89}{8}\right)\kappa_{S}-\frac{137}{384}\right]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\left[\left(\frac{89}{12}-\frac{89\nu}{6}\right)\kappa_{A}-\frac{71\delta}{192}+\frac{89\delta\kappa_{S}}{12}\right]+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\left[\frac{89\delta\kappa_{A}}{24}+\frac{71\nu}{192}+\left(\frac{89}{24}-\frac{89\nu}{12}\right)\kappa_{S}-\frac{71}{384}\right]$
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\left(\frac{89\delta\kappa_{A}}{24}+\frac{71\nu}{192}+\left(\frac{89}{24}-\frac{89\nu}{12}\right)\kappa_{S}-\frac{71}{384}\right)+\frac{833\nu^{2}}{216}-\frac{32537\nu}{1008}+\frac{26464169}{1524096}\,$, (10d)
@@
where the integrating factor is $e^{\int P(y)dy}$ and $C$ is a constant of integration that is fixed by enforcing that $R(y_{0})=R_{0}$. In the present case, $R\equiv e^{2}$. Since we require a solution that is in a PN series, we perform a Taylor series on the PN parameter in Eq. (13) to obtain the following solution for $e^{2}(y)$ till 2 PN and till $e_{0}^{2}$ accuracy
@@
where
$d_{\rm SS,circ}$ $=\left(-\frac{5}{496}+\frac{5\nu}{248}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}-\frac{5\delta}{248}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)+\left(-\frac{5}{496}+\frac{5\nu}{248}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}$
$+\left(\frac{5}{496}-\frac{5\nu}{248}\right)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}+\frac{5}{248}\delta\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\left(\frac{5}{496}-\frac{5\nu}{248}\right)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}$
$+\left(\frac{y_{0}}{y}\right)^{31/3}\left[\left(\frac{5}{496}-\frac{5\nu}{248}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}+\frac{5\delta}{248}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}+\left(\frac{5}{496}-\frac{5\nu}{248}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}\right$.
$\left.+\left(-\frac{5}{496}+\frac{5\nu}{248}\right)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}-\frac{5\delta}{248}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\left(-\frac{5}{496}+\frac{5\nu}{248}\right)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\right]\,$, (15a)
$d_{\rm NS,ecc}$ $=\left(\frac{y_{0}}{y}\right)^{19/3}+y^{2}\left[\left(\frac{y_{0}}{y}\right)^{25/3}\left(\frac{2833}{1008}-\frac{197\nu}{36}\right)+\left(\frac{y_{0}}{y}\right)^{19/3}\left(-\frac{2833}{1008}+\frac{197\nu}{36}\right)\right]+y^{3}\left[-\frac{377\pi}{72}\left(\frac{y_{0}}{y}\right)^{19/3}\right$.
$\left.+\frac{377\pi}{72}\left(\frac{y_{0}}{y}\right)^{28/3}\right]+y^{4}\biggl[\left(\frac{y_{0}}{y}\right)^{25/3}\left(-\frac{8025889}{1016064}+\frac{558101\nu}{18144}-\frac{38809\nu^{2}}{1296}\right)+\left(\frac{y_{0}}{y}\right)^{31/3}\left(-\frac{1193251}{3048192}-\frac{66317\nu}{9072}\right$.
$\left.+\frac{18155\nu^{2}}{1296}\right)+\left(\frac{y_{0}}{y}\right)^{19/3}\left(\frac{12635459}{1524096}-\frac{60781\nu}{2592}+\frac{10327\nu^{2}}{648}\right)\biggr]\,$, (15b)
$d_{\rm SO,ecc}$ $=\left(\frac{y_{0}}{y}\right)^{28/3}\left[\left(\frac{55\nu}{27}-\frac{157}{54}\right)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}-\frac{157\delta}{54}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right]+\left(\frac{y_{0}}{y}\right)^{19/3}\left[\frac{157\delta}{54}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}+\left(\frac{157}{54}-\frac{55\nu}{27}\right)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right]\,$, (15c)
$d_{\rm SS,ecc}$ $=\left(\frac{y_{0}}{y}\right)^{50/3}\biggl[\left(-\frac{725}{75392}+\frac{725\nu}{37696}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}-\frac{725\delta}{37696}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)+\left(-\frac{725}{75392}+\frac{725\nu}{37696}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}$
$+\left(\frac{725}{75392}-\frac{725\nu}{37696}\right)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}+\frac{725\delta}{37696}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\left(\frac{725}{75392}-\frac{725\nu}{37696}\right)\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\biggr]$

### v2

# Spin precession effects in the phasing formula of eccentric compact binary inspirals up to the second post-Newtonian order

@@ Abstract
Compact binary systems that emit gravitational waves (GW) are expected to exhibit orbital eccentricity together with generic spin configurations, leading to the precession of the orbital angular momentum, the individual spins, and the orbital plane around the total angular momentum. Although eccentric binaries with aligned spins have been extensively studied, closed-form post-Newtonian (PN) expressions that simultaneously capture both eccentricity and precessing-spin effects have remained unavailable. The presence of eccentricity significantly complicates orbital evolution, since solving the coupled differential equations generally requires numerical integration, thereby slowing waveform generation. In this work, we exploit the separation of timescales between the orbital motion, spin precession, and radiation reaction, and employ the precession-averaging method introduced by Morras et al. (2025) to effectively “average away” the explicit time dependence from the spin–orbit and spin–spin dynamics up to the second PN order. Building on this framework, we derive analytic phasing formulae from the evolution equations for the orbital frequency and eccentricity, treating eccentricity as a small parameter. Closed-form solutions for the eccentricity evolution and gravitational-wave phase are obtained up to the eighth power of the initial eccentricity $e_{0}$. The TaylorT2 approximant is generalized to incorporate spin-precession effects, and the orbital phase is calculated in both the time and frequency domains. To further enhance applicability, a resummation is performed on the TaylorT2 phasing formula to improve accuracy for moderate to high initial eccentricities. Taken together, these results provide efficient, closed-form phasing expressions that consistently capture the interplay of eccentricity and precessing spins, thereby enabling accurate and computationally tractable GW modeling for data analysis.
@@ I Introduction
A key part that comes after parameter estimation is the inference of formation channels of neutron stars and stellar-mass black holes. Eccentricity [21, 22, 23, 24, 25, 26, 27], spin-precession [28, 29, 30], and the combination of the two [31], among other observables, are extremely useful for tracing the formation channels of such compact objects [32]. It is to be noted that the absence of either eccentricity or spin-precession from template banks can lead to significant biases in parameter estimation [33, 34, 35, 36, 37, 38, 39, 40]. Also, since both introduce amplitude modulations in the GW signal, it becomes important to introduce both effects in waveform models in order to un-entangle the degeneracy caused by the presence of the two. Moreover, incorporating spin-precession effects into waveform models enables the extraction of additional binary parameters that are otherwise difficult to constrain in non-precessing systems. In such systems, these parameters tend to be strongly degenerate with others; precession mitigates this issue by causing the orbital plane to evolve in orientation during the inspiral, effectively providing multiple lines of sight to the same source [41]. Furthermore, inclusion of all known physical effects in waveform models will help in performing precision tests of General Relativity (GR) [42, 43, 44, 45, 46, 47, 48, 49].
So far, there has been significant progress in the development of eccentric waveform models. One of the first inspiral-merger-ringdown (IMR) models incorporating eccentricity was ENIGMA [50, 51], followed recently by its spinning counterpart ESIGMA [52]. PN inspired or inspiral only models include [53, 54, 55, 56, 57, 58], effective-one-body (EOB) models include [59, 60, 61, 62, 63], whereas NR-informed models include [64, 65, 66]. Similarly, for precessing spins, PN calculations include [67, 68, 69, 70, 71], EOB models include [72, 73, 74], Phenom models include [75, 76], and NR-informed models include [77]. Meanwhile, progress is beginning to be made in the combined modeling of precession and eccentricity [78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89].
@@
This paper is an update to both [53] and [101]. Compared to the aligned-spin eccentric phasing of [101], in this paper we generalize the phasing formula to precessing-spins in elliptical orbits. We give the phasing formula in both time and frequency domain up to 2 PN, and to an accuracy of the eighth power of the initial eccentricity $e_{0}$. We use inputs from [87] to calculate two Taylor approximants: TaylorT2 and TaylorF2 in the time and frequency domain respectively. We find that the phasing formulae are functions of intrinsic orbital parameters like mass ratios, spins, and quadrupole deformability factors. We also estimate the domain of validity of the phasing formula in both time and frequency domain by performing cycles computation and a mismatch analysis with existing waveform implementations on LALSuite[102, 103], an algorithm library for GW analysis.
@@
This paper is organized as follows: In Sec. II, we first present the inputs, that is, the evolution equations for the eccentricity weighted orbital frequency (or the PN parameter $y$) and the time-eccentricity $e_{t}$ (to be called as eccentricity henceforth), which we denote as $e$ in this paper. We then turn the evolution equations with respect to time as a single evolution equation for the eccentricity $e$ as a function of the PN parameter $y$ using chain rule. This particular equation can only be solved exactly at the Newtonian order for arbitrary eccentricities. However, for higher PN orders, the evolution equation simply becomes untractable by currently available methods. Nevertheless, as shown by [53], if one treats eccentricity as a small parameter, the differential equation (DE) can be solved exactly, and a closed-form solution can be found up to some order of the initial eccentricity $e_{0}$. However, the coefficient of each PN order in the eccentricity expanded DE was constant for the non-spinning and even in the aligned spins case. For generic spins, the coefficients themselves are time dependent, leading to further complications in solving the eccentricity evolution DE. To get around it, Refs. [81, 87] showed that the early inspiral binary problem could be treated as a multiple scale analysis (MSA) problem, which is possible since timescales for the orbit, spin-precession, and radiation-reaction can be clearly segregated. This allows the PN coefficients dependent on the spins to be precession-averaged, effectively taking away the time-dependence from the spin dynamics at the radiation-reaction timescale. Hence, these effective constants can now be put back into the eccentricity evolution DE and then solved as an ordinary DE, which we perform. The eccentricity solution as a function of the eccentricity-weighted orbital frequency $y$ is then utilized to find the time to coalescence $t$ as a function of $y$, and consequently, the time-domain orbital phasing $\phi$, or TaylorT2. We then perform a simple resummation of the TaylorT2 phasing. Finally, we use the Shifted Uniform Asymptotics (SUA) method introduced by [71] to find the frequency domain phasing $\Psi$.
@@
In Sec. III, we first calculate the number of cycles a real GW signal sweeps across a particular detector frequency range. We compare three different spin scenarios (almost aligned, almost anti-aligned, and perpendicular) for three different detector configurations and mass ranges. This we show for two different initial eccentricities. We then calculate the number of cycles swept by the numerically calculated TaylorT2 phase from the evolution equations and compare it with the analytical small-eccentricity expanded, as well as the resummed TaylorT2 phase number of cycles. Through this comparison, we show the domain of validity of our analytic TaylorT2 phase. We show that the $e_{0}^{8}$ expanded TaylorT2 phase incurs less than one cycle, compared to the numerically evaluated TaylorT2 phase, for initial eccentricity $e_{0}$ as high as 0.7. We also show that the resummed TaylorT2 is valid for initial eccentricities as high as 0.8. We then add our newly computed phasing correction to a frequency domain Taylor waveform and perform mismatch studies to ascertain the domain of validity of our PrecTaylorF2Ecc with respect to the eccentric circular aligned-spins waveform TaylorF2Ecc and the numerically computed eccentric spin-precessing waveform pyEFPE by [87].
@@
where $\bm{S}_{i}=m_{i}^{2}\bm{\chi}_{i}$, and $\bm{\chi}_{A/S}=\frac{\bm{\chi}_{1}\mp\bm{\chi}_{2}}{2}$. However, the approximants in [104] have been expressed in ($\bm{\chi}_{1},\bm{\chi}_{2},q_{1},q_{2}$) notation for ease of use for future calculations. We also use symmetric/anti-symmetric spin induced quadrupole moment constants as $\kappa_{A/S}=\frac{q_{1}\mp q_{2}}{2}$ in our results, where $q_{i}$ were defined in [87] after Eq. (C6). Furthermore, we replace all the individual masses $m_{i}$ with $\nu=\frac{m_{1}\,m_{2}}{M^{2}}$ and $\delta=\sqrt{1-4\nu}$, where $M$ is the total mass of the binary system.
@@ II.1 Eccentricity and PN parameter evolution equations
The evolution equations of $y$ and $e$ are given up to 2 PN for fully spinning and eccentric orbits by
@@
where $\mathcal{D}\equiv\frac{M}{\left(1-e^{2}\right)^{3/2}}\frac{d}{dt}$, $y=\frac{\left(M\omega\right)^{1/3}}{\sqrt{1-e^{2}}}$, and the non-spinning (NS) coefficients are given up to 2 PN by [87].
@@
The spin-orbit (SO) terms are given up to 2 PN in terms of $\chi_{eff}=\bm{\hat{L}}\cdot\bm{s}$ and $\delta\chi=\bm{\hat{L}}\cdot\left(\bm{s}_{1}-\bm{s}_{2}\right)$ by [87]
@@ II.2 Eccentricity evolution
Eqs. (2) can be combined to obtain the eccentricity evolution with respect to the PN parameter $y$. The DE cannot be solved for arbitrary eccentricities using current methods and hence needs to be expanded in small eccentricity in order to obtain an analytical solution. The $e^{2}$ expanded DE has the following form, after precession averaging of the various PN coefficients appearing in the evolution equations (2)
@@
$c_{\rm SS,circ}$ $=\left(\frac{5\nu}{12}-\frac{5}{48}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}-\frac{5\delta}{24}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)+\left(\frac{5}{48}-\frac{5\nu}{12}\right)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}$
$+\frac{5\delta}{24}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}-\frac{5}{48}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}+\frac{5}{48}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\,$, (10a)
$c_{\rm NS,ecc}$ $=-\frac{19}{3y}+y\left(\frac{197\nu}{18}-\frac{2833}{504}\right)-\frac{377\pi y^{2}}{24}+y^{3}\left(\frac{833\nu^{2}}{216}-\frac{32537\nu}{1008}+\frac{26464169}{1524096}\right)\,$, (10b)
$c_{\rm SO,ecc}$ $=\frac{157\delta}{18}\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}+\left(\frac{157}{18}-\frac{55\nu}{9}\right)\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\,$, (10c)
$c_{\rm SS,ecc}$ $=\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)\left[\left(\frac{89\nu}{2}-\frac{89}{4}\right)\kappa_{A}-\frac{137\delta}{192}-\frac{89\delta\kappa_{S}}{4}\right]$
$+\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}\left[-\frac{89\delta\kappa_{A}}{8}-\frac{21\nu}{4}+\left(\frac{89\nu}{4}-\frac{89}{8}\right)\kappa_{S}-\frac{137}{384}\right]$
$+\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}\left[-\frac{89\delta\kappa_{A}}{8}+\frac{641\nu}{96}+\left(\frac{89\nu}{4}-\frac{89}{8}\right)\kappa_{S}-\frac{137}{384}\right]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}\left[\left(\frac{89}{12}-\frac{89\nu}{6}\right)\kappa_{A}-\frac{71\delta}{192}+\frac{89\delta\kappa_{S}}{12}\right]$
$+\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\left[\frac{89\delta\kappa_{A}}{24}+\frac{261\nu}{32}+\left(\frac{89}{24}-\frac{89\nu}{12}\right)\kappa_{S}-\frac{71}{384}\right]$
$+\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\left[\frac{89\delta\kappa_{A}}{24}-\frac{89\nu}{12}+\left(\frac{89}{24}-\frac{89\nu}{12}\right)\kappa_{S}-\frac{71}{384}\right]\,$, (10d)
@@
where the integrating factor is $e^{\int P(y)dy}$ and $C$ is a constant of integration that is fixed by enforcing that $R(y_{0})=R_{0}$. In the present case, $R\equiv e^{2}$. Since we require a solution that is in a PN series, we perform a Taylor series on the PN parameter in Eq. (13) to obtain the following solution for $e^{2}(y)$ up to 2 PN and to $e_{0}^{2}$ accuracy
@@
where $d_{\rm NS,ecc}$ is provided in [105] and $d_{\rm SO,ecc}$ is provided in [101]. The precessing coefficients are as follows-
$d_{\rm SS,circ}$ $=\left(-\frac{5}{496}+\frac{5\nu}{124}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}-\frac{5\delta}{248}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)$
$\quad-\frac{5}{496}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}+\left(\frac{5}{496}-\frac{5\nu}{124}\right)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}$
$\quad+\frac{5}{248}\delta\,\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}+\frac{5}{496}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}$
$\quad+\left(\frac{y_{0}}{y}\right)^{31/3}\left[\left(\frac{5}{496}-\frac{5\nu}{124}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)^{2}+\frac{5\delta}{248}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)\right$.
$\quad\left.+\frac{5}{496}\left(\bm{\hat{\bm{L}}}\cdot\bm{\chi}_{S,0}\right)^{2}+\left(-\frac{5}{496}+\frac{5\nu}{124}\right)\bm{\chi}_{A,0}\cdot\bm{\chi}_{A,0}\right$.
$\quad\left.-\frac{5\delta}{248}\bm{\chi}_{A,0}\cdot\bm{\chi}_{S,0}-\frac{5}{496}\bm{\chi}_{S,0}\cdot\bm{\chi}_{S,0}\right]\,$, (15a)
$d_{\rm SS,ecc}$ $=\left(\frac{y_{0}}{y}\right)^{50/3}\biggl[\left(-\frac{725}{75392}+\frac{725\nu}{18848}\right)\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\right)^{2}-\frac{725\delta}{37696}\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{A,0}\right)\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\right)$
$\quad-\frac{725}{75392}\left(\bm{\hat{\bm{L}}}\!\cdot\!\bm{\chi}_{S,0}\right)^{2}+\left(\frac{725}{75392}-\frac{725\nu}{18848}\right)\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{A,0}$
$\quad+\frac{725\delta}{37696}\bm{\chi}_{A,0}\!\cdot\!\bm{\chi}_{S,0}+\frac{725}{75392}\bm{\chi}_{S,0}\!\cdot\!\bm{\chi}_{S,0}\biggr]$